solve the equation by completing the square.
(3x)square -6x-1 = 0

Answer by Nate(3500)

3x^2 - 6x - 1 = 0
3x^2 - 6x = 1
x^2 - 2x = 1/3
(x - 1)^2 = 4/3
x - 1 = +- 2sqrt(3)/3
x = 1 +- 2sqrt(3)/3